Evaluate. pleae help!!
Problem #1 |4| + |-3|



Problem #2 |-9| - |-4|

Answer by nabla(475)

Absolute value means take the magnitude of the number--- the size--- irrespective of the sign.
|4| + |-3| =4+3=7

|-9| - |-4|=9-4=5